OpenStudy (anonymous):

In a triangle, the ratio of the measures of the sides is 2:3:4 and the perimeter is 72. Find the measures of the sides of the triangle.

hero (hero):

2x + 3x + 4x = 72 9x = 72 x = 72/9 x = 8 2x = 2(8) = 16 3x = 3(8) = 24 4x = 4(8) = 32 16 + 24 + 32 = 72

OpenStudy (anonymous):

one side measures 16. another side measures 24. the third side measures 32